The bicycle industry consists of seven firms. Firms 1, 2, 3, 4 each has 10% market share, and firms 5,6,7 each has 20% market share. Using the concentration measures, answer the following questions:

(i) Calculate 4-firm concentration ratio for this industry.

(ii) Calculate the Herfindahl-Hirschman index (HHI) for this industry.

(iii) Now, suppose that firms 1 and 2 merge, so that the new firm will have a market share of 20%.

1) Calculate the post merger I(x)

2) Calculate the change in the I(x) caused by the merger.

3) Will this merger be challenged in the U.S.? In Canada?
Answer #1

Answer:

A) 4 - firm concentration ratio= sum of market share of top 4 firms=20+20+20+10=70%

B)HHI =〉 (marketshare)

HHI=10^2+10^2+10^2+10^2+20^2+20^2+20^2=1600

C)

1) After merger

HHI=20^2+10^2+10^2+20^2+20^2+20^2=1800

2) Change in HHI=1800-1600=200

3)Since market was already moderately concentrated with 1600 HHI and after merger HHI is increased by 200 so it is assumed to increase the market power so US and Canada won't allow this merger.